### v1

Priority-Neutral Matching Lattices Are Not Distributivethanks: We thank Ata Atay and Nicole Immorlica for helpful conversations.
Clayton Thomas Microsoft Research. — E-mail: clathomas@microsoft.com.
(April 2, 2024)
Stable matchings are a cornerstone of market design, with numerous practical deployments backed by a rich, theoretically-tractable structure. However, in school-choice problems, stable matchings are not Pareto optimal for the students. Priority-neutral matchings, introduced by [Ren22], generalizes the set of stable matchings by allowing for certain priority violations, and there is always a Pareto-optimal priority-neutral matching. Moreover, like stable matchings, the set of priority-neutral matchings forms a lattice [Ren22].
We study the structure of the priority-neutral lattice. Unfortunately, we show that much of the simplicity of the stable matching lattice does not hold for the priority-neutral lattice. In particular, we show that the priority-neutral lattice need not be distributive. Moreover, we show that the greatest lower bound of two matchings in the priority-neutral lattice need not be their student-by-student minimum, answering an open question of [Ren22]. This show that many widely-used properties of stable matchings fail for priority-neutral matchings; in particular, the set of priority-neutral matchings cannot be represented by via a partial ordering on a set of rotations. However, by proving a novel structural property of the set of priority-neutral matchings, we also show that not every lattice arises as a priority-neutral lattice, which suggests that the exact nature of the family of priority-neutral lattices may be subtle.
1 Introduction
Stable matchings enjoy a celebrated position within both the theory and practice of market design. A vast theoretical literature establish rich structural properties of stable matchings, and widespread practical deployments harness this theory to solve real-world problems.
However, in some settings such as students-schools assignment, enforcing stability may be overly stringent, in two ways. First, stability may be overly stringent as a requirement. This is because students often have little ability to match outside of their centrally-assigned school; thus, priority violations—i.e., pairs $(d,h)$ where $d$ prefers $h$ to their match, and $h$ admits a student of lower priority than $d$ (or has an empty seat)—may be undesirable due to a sense of unfairness, but may not compromise the integrity of the matching market overall. Second, stability may be overly stringent in its implications. Namely, stability implies forgoing Pareto-optimality for the students. For instance, [APR09] report that in 2006 in New York City, over 4,000 grade 8 students could be made better off without worsening the match of any student. This leads to an important question: is there a natural alternative to stability which permits Pareto-optimal matches, while paying due respect to priority violations?111 The classical approach to Pareto-optimal matching in the school-choice problem is the Top Trading Cycles (TTC) mechanism [SS74, AS03]. While TTC is often a desirable mechanism, it (informally speaking) ignores priority violations completely, motivating the search for solution concepts that make a more integrated compromise between stability and Pareto-optimality.
Motivated by this question, [Ren22] proposes a generalization of stability, termed priority-neutrality. A matching $\mu$ is priority-neutral if, for each student $d$ whose priority is violated by $\mu$, it is not possible for another matching $\nu$ to make $d$ better off, unless $\nu$ also violates the priority of some student who is made worse off. Since stable matchings violate no student’s priority, each stable matching is priority-neutral. Additionally, priority-neutral matchings also allow for certain violations of $d$’s priority when, in some sense, $d$ cannot be improved without undue effects on others. [Ren22] proves that the set of priority-neutral matchings has mathematical structure: like the set of stable matchings, the set of priority-neutral matchings forms a lattice under the preferences of the students.222 See Section 2 for the definition of the ordering relation on matchings, and the definition of a lattice. Thus there is a unique student-optimal priority-neutral matching, which [Ren22] additionally shows is Pareto-efficient.333 [Ren22] is far from the first work to pursue the agenda of Pareto-improving stable matchings; see Section 1.1. Perhaps most notably, [EM20] introduce legal matchings, were a set $\mathsf{L}$ of matchings is legal if $\mathsf{L}=\{\mu\ \allowbreak|\ \allowbreak\text{$\forall\nu\in\mathsf{L}:$ \allowbreak no student $d$'s priority is violated by $\mu$ at school $\nu(d)$}\}$. [EM20] prove there is a unique legal set, and these legal matchings also form a lattice with a unique Pareto-efficient student-optimal outcome. However, the self-referential nature of the definition of a legal set motivates the search for other notions such as priority-neutrality, which [Ren22] posits may be a simpler definition than legality. Thus, known results suggest that priority-neutral matchings may be as theoretically tractable as stable matchings.
In this paper, we show that a crucial property of the set of stable matchings does not in fact carry over to priority-neutral matchings. In particular, we show that the priority-neutral lattice need not be distributive.444 A distributive lattice is one in which the equation $\mu\wedge(\nu\vee\xi)=(\mu\wedge\nu)\vee(\mu\wedge\xi)$ always holds, where $\wedge$ denotes greatest lower bounds and $\vee$ denotes least upper bounds. This lattice-theoretic property, while perhaps initially seemingly abstract, turns out to be intimately related to the theory of stable matchings. For instance, distributivity exactly characterizes stable matching lattices; i.e. a lattice arises as the set of stable matchings if and only if it is distributive [Knu74, Bla84, IL86]. Moreover, distributivity underpins the classical representation of the stable matching lattice by a partially ordered set of rotations [GI89] (since Birkhoff’s representation theorem [Bir37] says that distributive lattices are exactly those which can be represented as the downward-closed subsets of some partial order). This representation is widely exploited in both algorithmic and theoretical works on stable matchings (see, e.g., [Man13, EIV23]). Our result proves that the priority-neutral lattice cannot be represented by a partial order on some set of rotations.
Additionally, our construction resolves an open question of [Ren22]: we show that—unlike for stable matchings—the greatest lower bound of two matchings in the priority-neutral lattice need not be their student-by-student minimum. This result may be surprising: in contrast to stable matching lattices, as well as other well-known lattices from economic theory such as the lattice of Walrasian price vectors in an economy with gross substitutes consumer [GS99], lattice operations for priority-neutral matchings are not given by coordinate-by-coordinate minima / maxima.
This fact raises a natural question: what is the family of lattices that can arise as the set of priority-neutral matchings? We make progress towards this answer by proving that some lattices cannot be represented in this way, i.e., we show there exists a lattice $Q$ such that no set of preferences (and priorities) exist for which $Q$ is the priority-neutral lattice. Informally, our proof rules out certain lattices by showing that if there are many distinct priority-neutral matchings near the bottom (i.e., student-pessimal) priority-neutral matching, then there must also be many distinct priority-neutral matchings further above these elements. To complete the argument, we generalize the techniques of [Ren22] and exploit properties of other related matching lattices [EM20, FZ22] to prove a novel structural property of the priority-neutral lattice; for details, see Section 4.
In summary, we prove two main results. First, priority-neutral lattices fail to satisfy distributivity, a crucially important structural property of many lattices of economic importance. Second, not all lattices arise as priority-neutral lattices, showing that the precise nature of priority-neutral lattices may be subtle. Future works may give additional structural results clarifying how the priority-neutral lattice behaves. They may also look for alternative definitions which also reach efficient and fair matchings, according to appropriate new notions, with a more tractable structure.
1.1 Related Work
This work primarily relates to two strands of literature: papers studying the structure of the set of stable matchings, and papers studying Pareto-improvements of stable matchings. Works on stable matchings follow from the seminal [GS62]. For structural properties specifically, [GI89] is the classical reference; see also [EIV23, Section 1.4]. Key individual works include [Knu74, Bla84, IL86], which establish that stable matching lattices are exactly characterized by distributive lattices. Other works establishing fundamental properties of stable matching mechanisms include [GS85a, MW71, Rot82, Rot84, DF81, GS85b].
[Kes10] introduces the $\mathsf{EADAM}$ mechanism, which has become a centerpoint of the literature that seeks Pareto-improvements to stable matchings. [TY14] provides a simpler algorithm for computing $\mathsf{EADAM}$. [Ren22] introduces priority-neutral matchings, proved they form a lattice, and proved that the student-optimal matching in this lattice equals the outcome of $\mathsf{EADAM}$.
Other than priority-neutral matchings, likely the most significant prior theoretical framework for Pareto-improving the set of stable matchings is the set of legal matchings [EM20], which we denote by $\mathsf{L}$; see footnote 3 for a definition. [Ren22] proves that every priority-neutral matching is legal (and that there can be legal, but not priority-neutral, matchings). [EM20] proves that $\mathsf{L}$ forms a lattice, and the student-optimal matching in $\mathsf{L}$ equals the outcome of $\mathsf{EADAM}$. Thus, these two matching lattices are closely related: the priority-neutral lattice is contained within the legal lattice, and they have the same student-optimal element. However, [EM20] show in fact that $\mathsf{L}$ forms a distributive lattice which is closed under taking the student-by-student minimum and maximum; in contrast, our main result demonstrates that the structure of priority-neutral lattice can be quite different from that of the legal lattice, since priority-neutral lattices need not satisfy these fundamental properties. Nevertheless, we exploit the connection between legal and priority-neutral matchings, along with additional properties of the legal lattice proven in [FZ22], in our structural result in Section 4.
[EM20] follows the general idea of “stable sets” as in [NM44]; other papers taking similar or related approaches include [Ehl07, MVV11, AMV22, DE23], among others. [TDK20] define essentially stable matchings, which allow for Pareto-optimal matchings, but does not form a lattice (or characterize a unique Pareto-optimal matchings). [ACP+20, KS20, TZ21] compare mechanisms by the sets of priorities they violate. [CM23] studies a generalization of strategyproofness for $\mathsf{EADAM}$. [RTZ23] gives a novel characterization of $\mathsf{EADAM}$, as do many of the previously-mentioned papers.
2 Preliminaries
2.1 Stable matchings, priority violations, lattices, EADAM, and rotations
Stable matchings and priority violations.
We study matchings between applicants and institutions (outside of the exposition in Section 1, where we referred to students and schools). Denote the set of applicants by $\mathcal{A}$ and institutions by $\mathcal{I}$. For simplicity we assume that all matchings are one-to-one, i.e. that institutions have capacity $1$. We typically denote matchings by $\mu,\nu,\xi$, etc., applicants by $d$ (mnemonic: doctor), and institutions by $h$ (mnemonic: hospital); let $\mu(d)$ denote the match of $d$ in $\mu$ (possibly equal to $\emptyset$ if $d$ is unmatched). Each agent has strict preferences over the other side. We denote the preferences of applicant $d$ as $\succ_{d}$, and write $h_{1}\succ_{d}h_{2}$ if $d$ prefers institution $h_{1}$ to institution $h_{2}$. Applicants may have partial preference lists, indicating that they prefer to remain unmatched over being matched to any institution not on their list. The “preferences” of the institutions, also denoted by $\succ_{h}$ for an institution $h$, are referred to as priorities. This reflects the fact that we think of the institutions as “objects to be consumed” without preferences of their own.555 For example, in student-school matching, priorities of schools might be determined by pre-specified policies based on a student’s socio-economic status or distance from the school. For simplicity, we assume that each institution ranks all applicants in their priority lists.
We think of stability in terms of avoiding violating applicants’ priorities, as follows:
Definition 2.1.
For an applicant $d$, institution $h$, and matching $\mu$, we say that $d$’s priority at $h$ is violated in $\mu$ if $h\succ_{d}\mu(d)$ and $d\succ_{h}\mu(h)$. We say that $d$’s priority is violated in $\mu$ if there exists such an $h$. We denote by $\operatorname*{vio}(\mu)\subseteq\mathcal{A}$ the set of applicants whose priority is violated in $\mu$.
A matching $\mu$ is stable if no applicant’s priority is violated in $\mu$, i.e., if $\operatorname*{vio}(\mu)=\emptyset$.
The lattice of stable matchings.
A lattice under a partial order $\leq$ is a set $Q$ such that every pair of elements in $Q$ have a greatest lower bound and a least upper bound in $Q$. The greatest lower bound of $x,y\in Q$, also called their meet in $Q$ and denoted $x\wedge_{Q}y$, is the element $z\in Q$ such that $z\leq x$ and $z\leq y$, and moreover for each $w\in Q$ such that $w\leq x$ and $w\leq y$, we have $w\leq z$. Likewise, the least upper bound of $x,y\in Q$, also called their join in $Q$ and denoted $x\vee_{Q}y$, is the element $z\in Q$ such that $z\geq x$ and $z\geq y$, and for each $w\in Q$ such that $w\geq x$ and $w\geq y$, we have $w\geq z$. Since we study different lattices on sets of matchings, we often emphasize the relevant lattice $Q$ using the notation $\wedge_{Q}$ and $\vee_{Q}$. A lattice $Q$ is distributive if, for each $x,y,z\in Q$, we have $\mu\wedge_{Q}(\nu\vee_{Q}\xi)=(\mu\wedge_{Q}\nu)\vee_{Q}(\mu\wedge_{Q}\xi)$.
We now explain the classical lattice structure of stable matchings. We say that matching $\mu$ (weakly) dominates matching $\nu$, denoted $\mu\geq\nu$, if for each applicant $d$ we have $\mu(d)\succeq_{d}\nu(d)$. We write $\mu>\nu$ if $\mu\geq\nu$ and $\mu\neq\nu$. For two matchings $\mu,\nu$, let the coordinatewise maximum of $\mu$ and $\nu$ be the function $\xi$ from applicants to institutions given by $\xi(d)=\max_{\succ_{d}}\{\mu(d),\nu(d)\}$. Analogously, let their coordinatewise minimum of $\mu$ and $\nu$ be such that $\xi(d)=\min_{\succ_{d}}\{\mu(d),\nu(d)\}$. Note that for general $\mu$ and $\nu$, the coordinatewise maximum/minimum may not be a matching.
Theorem 2.2 (See, e.g., [GI89]).
The set of stable matchings is always nonempty, and forms a lattice under the dominance ordering. In this lattice, the least upper bound of any two stable matchings is their student-by-student maximum, and the greatest lower bound is analogously their student-by-student minimum. Moreover, a lattice $Q$ is isomorphic to a lattice of stable matchings for some set of preferences and priorities if and only if $Q$ is distributive.
We refer to the applicant-optimal stable matching as the applicant-proposing deferred acceptance (denoted $\mathsf{APDA}$) matching. Likewise, we term the institution-optimal version institutions-proposing deferred acceptance (denoted $\mathsf{IPDA}$).
Pareto improving stable matchings and EADAM.
A matching $\mu$ is Pareto-optimal if there does not exist a $\nu\neq\mu$ with $\nu\geq\mu$. The literature on Pareto-improving stable matchings has identified a central mechanism which is Pareto-optimal and dominates $\mathsf{DA}$, namely, [Kes10]’s efficiency-adjusted DA mechanism ($\mathsf{EADAM}$). One approach to defining $\mathsf{EADAM}$, pioneered by [TY14], is as follows:
Definition 2.3.
For a profile of preferences and priorities $P$, define the Kesten-Tang-Yu sequence as follows. Let $\mu_{1}$ denote the outcome of $\mathsf{APDA}(P)$ under preference and priorities $P=P_{1}$. Call an institution $h$ under-demanded in round $i$ if $h$ never rejects any applicant during the run of $\mathsf{APDA}(P_{i})$ (additionally, the “institution” $\emptyset$ representing going unmatched is always considered under-demanded). For each $i\geq 2$, let $P_{i}$ denote modifying the profile of preferences as follows: for each student $d$ matched in $\mu_{i-1}$ to an institution $h$ which is under-demanded in round $i-1$, remove all institution strictly before $h$ in $d$’s preference list. There must exist an $N$ such that $\mu_{n}=\mu_{m}$ for each $n,m>N$. The outcome of efficiency-adjusted DA ($\mathsf{EADAM}$) is $\mu_{N}$.
[TY14] prove that at least one institution is under-demanded in each round, and thus $\mathsf{EADAM}$ can be computed via at most $N$ runs of $\mathsf{APDA}$, where $N$ is the number of institutions. Moreover, $\mathsf{EADAM}$ is Pareto-optimal and dominates all stable matchings.
$\mathsf{EADAM}$ is closely related to priority-neutral matchings and will feature into our analysis. The same is true of another prior notion, namely the set of legal matchings [EM20], but we defer defining legal matchings and their properties (e.g., from [FZ22]) until they are needed in Section 4.
Rotations.
Throughout the paper, we use the notion of rotations [GI89] to describe different matchings. A rotation is an ordered list of agents $\alpha=[(h_{1},d_{1}),\allowbreak(h_{2},d_{2}),\allowbreak\ldots,\allowbreak(h_{k},d_{k})]$, where each $h_{i}\in\mathcal{I}$ and $d_{i}\in\mathcal{A}$. If $\mu$ is a matching where $d_{i}$ is paired to $h_{i}$ for each $i=1,\ldots,k$, then we let $\alpha\mu$ denote the matching which differs from $\mu$ only by matching $d_{i}$ to $h_{i+1}$ for $i=1,\ldots,k$ modulo $k$. (The action of a rotation can be visualized as moving the parentheses in $\alpha$ over one place, modulo $k$.) Matching $\alpha\mu$ is also called the elimination of rotation $\alpha$ from $\mu$. In Section 3, we use rotations solely as notation, i.e., one can treat them as names for different matching. In Section 4, we will exploit many properties of rotations from the literature as needed; see [GI89] or [EIV23] for a treatment.666 Note that we will use the convention that $\alpha\mu\geq\mu$, i.e. all applicants weakly prefer $\alpha\mu$ to $\mu$. In contrast to this, [GI89] (resp., [EIV23]) define rotations such that $\mu\geq\alpha\mu$, i.e., such that all men (resp., workers) weakly prefer $\mu$ to $\alpha\mu$. Other than that this difference in convention, the notions are identical. We write $\alpha_{1}\alpha_{2}\ldots\alpha_{\ell}\mu$ to mean $\alpha_{1}(\alpha_{2}(\ldots(\alpha_{\ell}\mu)\ldots))$; when the “base” matching $\mu$ is understood, we also simply denote this by $\alpha_{1}\alpha_{2}\ldots\alpha_{\ell}$.
2.2 Priority-neutral matchings and their join and meet operations
Priority-neutrality.
We now turn to the central notion of this paper, priority-neutral matchings [Ren22]. In words, a matching $\mu$ is priority-neutral if it’s impossible to make any student whose priority is violated in $\mu$ better off, without violating the priority of some student who is made worse off. To fully spell out this definition, we introduce the terminology “priority-correcting adjustment” to aid in discussing matchings which are not priority-neutral. Namely, we have:
Definition 2.4.
For matchings $\xi$ and $\nu$, we say that $\nu$ is a priority-correcting adjustment of $\xi$ if:
  1. 1.
    There exists an applicant $d_{*}\in\operatorname*{vio}(\xi)$ with $\nu(d_{*})\succ_{d_{*}}\xi(d_{*})$.
  2. 2.
    For each $d$ such that $d\in\operatorname*{vio}(\nu)$, we have $\nu(d)\succeq_{d}\xi(d)$.
A matching $\mu$ is priority-neutral if there does not exist any priority-correcting adjustment of $\mu$. We denote the set of priority-neutral matchings by $\mathsf{PN}$.
$\mathsf{PN}$ generalizes the set of stable matchings, and [Ren22] proves that $\mathsf{PN}$ also forms a lattice. Moreover, [Ren22] proves that the set $\mathsf{PN}$ is closed under the coordinatewise maximum operation, so the join operation in the $\mathsf{PN}$ lattice is also given by the coordinatewise maximum. Formally:
Theorem 2.5 ([Ren22]).
For any profile of priorities and preferences, $\mathsf{PN}$ is a lattice under $\leq$, and this lattice contains every stable matching. Additionally, the student-pessimal element of $\mathsf{PN}$ is the institution-optimal stable matching, and the student-optimal element of $\mathsf{PN}$ equals the outcome of $\mathsf{EADAM}$. Finally, for any two matchings $\mu,\nu$ in $\mathsf{PN}$, the least upper bound of $\mu$ and $\nu$ in $\mathsf{PN}$ equals their coordinatewise maximum.
A simple example showing greatest lower bounds need not be the coordinatewise minimum.
Perhaps surprisingly, we will show in this paper that the greatest lower bound between two matchings in $\mathsf{PN}$ may not be their coordinatewise minimum. Before proceeding, we first describe how this type of situation is possible using a simple example.
Example 2.6.
Consider the following three matchings:
$\mu_{0}$ =$=\{(h_{1},d_{1}),(h_{2},d_{2}),(h_{3},d_{3}),(h_{4},d_{4}),(h_{5},d_{5}),(h_{6},d_{6})\}$
$\mu_{L}$ =$=\{(h_{1},d_{2}),(h_{2},d_{1}),(h_{3},d_{4}),(h_{4},d_{3}),(h_{5},d_{5}),(h_{6},d_{6})\}$
$\mu_{R}$ =$=\{(h_{1},d_{1}),(h_{2},d_{2}),(h_{3},d_{4}),(h_{4},d_{3}),(h_{5},d_{6}),(h_{6},d_{5})\}$
In words, relative to $\mu_{0}$, matching $\mu_{L}$ separately swaps the matches of $d_{1},d_{2}$ and (separately) of $d_{3},d_{4}$, and $\mu_{R}$ swaps $d_{3},d_{4}$ and $d_{5},d_{6}$. Let the applicant’s preferences be such $\mu_{L}$ and $\mu_{R}$ are both preferred to $\mu_{0}$ by every applicant who changes match between them. Then, within the set of matchings $Q=\{\mu_{0},\mu_{L},\mu_{R}\}$, matchings $\mu_{L}$ and $\mu_{R}$ have a greatest lower bound $\mu_{L}\wedge_{Q}\mu_{R}=\mu_{0}$, which is different from their coordinatewise minimum $[(h_{1},d_{1}),(h_{2},d_{2}),(h_{3},d_{4}),(h_{4},d_{3}),(h_{5},d_{5}),(h_{6},d_{6})]$.
Notation for joins and meets.
In Section 3, we embed an example like 2.6 into $\mathsf{PN}$. In view of such examples, we introduce separate notation for the coordinatewise minimum and the greatest lower bound of two matchings in $\mathsf{PN}$. Namely, let $\mu\wedge_{\mathsf{L}}\nu$ denote the coordinatewise minimum of $\mu$ and $\nu$,777 This notation is inspired by the legal set of matchings $\mathsf{L}$ from [EM20]; as we will exploit in Section 4, coordinatewise minimums $\wedge_{\mathsf{L}}$ give greatest lower bounds in the legal lattice $\mathsf{L}$. and let $\mu\wedge_{\mathsf{PN}}\nu$ denote their greatest lower bound in $\mathsf{PN}$. Since by Theorem 2.5, least upper bounds equal coordinatewise maximums in $\mathsf{PN}$ (and since this is also true in all other matching lattices we consider), we use the notation $\mu\vee\nu$ for least upper bounds (i.e., this always denotes the coordinatewise maximum).
3 Priority-Neutral Lattices Are Not Distributive
Recall that, like the classical set of stable matchings, the set of priority-neutral matchings has the mathematical structure of a lattice [Ren22]. In this section, we show that unfortunately, the $\mathsf{PN}$ lattice actually lacks much of the structure the lattice of stable matchings has. We construct a set of preferences and priorities such that $\mathsf{PN}$ is not be distributive, and that greatest lower bounds in $\mathsf{PN}$ are not given by coordinatewise minimums.
Figure 1 shows our construction. Figure 1 gives the priorities and preferences of the construction. We will show that Figure 1 displays the entire lattice of priority-neutral matchings for this instance, where $\mu\leq\nu$ if any only if there is a path from $\mu$ to $\nu$ consisting of edges ordered in their upward direction. The matchings in Figure 1 are described in terms of the rotations given in Figure 1, as discussed in Section 2.1. For brevity, we omit the “base” matching $\mu_{0}$ from our notation for different matchings; each time we write $\alpha_{1}\ldots\alpha_{k}$ through this section for rotations $\alpha_{1},\ldots,\alpha_{k}$, we mean the matching $\alpha_{1}\ldots\alpha_{k}\mu_{0}$.888 For example, the matching denoted $\alpha$ equals $[(h_{A},d_{B}),(h_{B},d_{A}),(h_{C},d_{C}),(h_{D},d_{D}),(h_{E},d_{E}),(h_{F},d_{F}),(h_{Z},d_{Z})]$, and the matching $\beta\gamma$ equals $[(h_{A},d_{A}),(h_{B},d_{B}),(h_{C},d_{D}),(h_{D},d_{C}),(h_{E},d_{F}),(h_{F},d_{E}),(h_{Z},d_{Z})]$.
Institutions’ Priorities:
$h_{A}$ $h_{B}$ $h_{C}$ $h_{D}$ $h_{E}$ $h_{F}$ $h_{Z}$
$d_{A}$ $d_{B}$ $d_{C}$ $d_{D}$ $d_{E}$ $d_{F}$ $d_{Z}$
$d_{B}$ $d_{A}$ $d_{Z}$ $d_{C}$ $d_{F}$ $d_{E}$ $d_{A}$
$d_{E}$ $d_{D}$ $d_{Z}$
Applicants’ Preferences:
$d_{A}$ $d_{B}$ $d_{C}$ $d_{D}$ $d_{E}$ $d_{F}$ $d_{Z}$
$h_{B}$ $h_{A}$ $h_{D}$ $h_{C}$ $h_{F}$ $h_{E}$ $h_{E}$
$h_{Z}$ $h_{B}$ $h_{C}$ $h_{D}$ $h_{A}$ $h_{F}$ $h_{C}$
$h_{A}$ $h_{E}$ $h_{Z}$
(a) Preferences and Priorities.
(b) Priority-neutral lattice.
Institution-optimal stable matching $\mu_{0}$:
$h_{A}$ $h_{B}$ $h_{C}$ $h_{D}$ $h_{E}$ $h_{F}$ $h_{Z}$
$d_{A}$ $d_{B}$ $d_{C}$ $d_{D}$ $d_{E}$ $d_{F}$ $d_{Z}$
Relevant rotations:
$\alpha$ $=[(h_{A},d_{A}),(h_{B},d_{B})]$
$\beta$ $=[(h_{C},d_{C}),(h_{D},d_{D})]$
$\gamma$ $=[(h_{E},d_{E}),(h_{F},d_{F})]$
(c) Matching and rotations used to denote the matchings in Figure 1.
Figure 1: Set of priorities and preferences witnessing Theorem 3.2, along with the priority-neutral lattice (with the stable matchings shaded in green).
We start by giving an high-level description of the important properties of the example in Figure 1. The crucial fact about this construction is that $\alpha\beta$ and $\beta\gamma$ are priority-neutral, but $\beta$ is not priority-neutral. This will immediately imply that $\mathsf{PN}$ is not closed under coordinatewise minimums; indeed, $\alpha\beta$ and $\beta\gamma$ serve a role analogous to $\mu_{L}$ and $\mu_{R}$ in 2.6. We found a construction with this property by combining two things. First, there is an applicant $d_{Z}$ can be Pareto-improved in $\mu_{0}$, the institution-optimal stable matching; however, either of the rotations $\alpha$ or $\gamma$ make this Pareto-improvement no longer possible.999 The most involved part of the proof below involves arguing that in $\alpha\beta$ and $\beta\gamma$, not only is $d_{Z}$ not Pareto-improvable, but $d_{Z}$ has no priority-correcting adjustment (2.4). Second, $d_{Z}$’s priority is violated in $\beta$, and additionally $d_{Z}$ can Pareto-improve in $\beta$ (as in $\mu_{0}$), showing that $\beta\notin\mathsf{PN}$. As we argue below, all of the other relevant combinations of rotations will be priority-neutral matchings, showing that $\mathsf{PN}$ is given by the lattice in Figure 1.
Proposition 3.1.
For the preferences and priorities given in Figure 1, priority-neutral lattice $\mathsf{PN}$ is given by Figure 1.
Proof.
To begin, we check that the stable matching lattice, which is always contained in $\mathsf{PN}$ by Theorem 2.5, is given by the green-shaded matchings in Figure 1. First, the institution-optimal matching equals $\mu_{0}$, as defined in Figure 1, since all institutions have distinct top-priority applicants equal to their match in $\mu_{0}$. Second, one can run applicant-proposing deferred acceptance to see that $\alpha\gamma$ is the applicant-optimal stable matching. Third, one can verify directly that the matchings $\alpha$ and $\gamma$ have no priority violations, and are thus stable; since no other distinct matchings $\mu^{\prime}$ exist with $\mu_{0}<\mu^{\prime}<\alpha\gamma$, these green shaded matchings are all stable matchings.101010 In full detail, in $\alpha$, all institutions except for $h_{A}$ and $h_{B}$ are matched to their top-priority applicants; while $d_{E}$ prefers $h_{A}$ to their match in $\alpha$, we have $d_{B}\succ_{h_{A}}d_{E}$, so $\alpha$ is stable. Similarly, in $\gamma$, all institutions except for $h_{E}$ and $h_{F}$ are matched to their top-priority applicants; while $d_{Z}$ prefers $h_{E}$ to their match in $\gamma$, we have $d_{F}\succ_{h_{E}}d_{Z}$, and $\gamma$ is stable. Alternatively, the methods of [GI89] can be used to directly compute the set of rotation representing the stable matching lattice (which are $\alpha,\gamma$ with no predecessor relations).
Next, we verify that the applicant-optimal priority-neutral matching is $\alpha\beta\gamma$. By Theorem 2.5, this matching equals the $\mathsf{EADAM}$ outcome, which can be shown to equal $\alpha\beta\gamma$ by the procedure in 2.3. (Indeed, in the second round of the Kesten-Tang-Yu sequence, $d_{Z}$ will be fixed at $h_{Z}$, and $d_{C}$ and $d_{D}$ will match to their top choices $h_{D}$ and $h_{C}$, so $\alpha\beta\gamma$ equals $\mathsf{EADAM}$ and this matching is the highest element of the $\mathsf{PN}$ lattice.)
Now, three relevant matchings remain which we have not yet considered: $\beta$, $\alpha\beta$, and $\beta\gamma$. Since these are the only not-yet-considered matchings $\mu$ such that $\mu_{0}<\mu<\alpha\beta\gamma$, we know that all non-stable matchings in $\mathsf{PN}$ must be in $\{\beta,\alpha\beta,\beta\gamma\}$ by Theorem 2.5. We will proceed by showing that $\beta\notin\mathsf{PN}$, but $\alpha\beta\in\mathsf{PN}$ and $\beta\gamma\in\mathsf{PN}$.
We now show $\beta\notin\mathsf{PN}$. To see this, consider $\beta$, and define the matching $\nu$ as follows:
$\beta:$ $h_{A}$ $h_{B}$ $h_{C}$ $h_{D}$ $h_{E}$ $h_{F}$ $h_{Z}$ $d_{A}$ $d_{B}$ $d_{D}$ $d_{C}$ $d_{E}$ $d_{F}$ $d_{Z}$     $\nu_{*}:$ $h_{A}$ $h_{B}$ $h_{C}$ $h_{D}$ $h_{E}$ $h_{F}$ $h_{Z}$ $d_{E}$ $d_{B}$ $d_{D}$ $d_{C}$ $d_{Z}$ $d_{F}$ $d_{A}$
We claim that $\nu_{*}$ is a priority-correcting adjustment of $\beta$ for $d_{Z}$. To see this, note that in $\beta$, applicant $d_{Z}$’s priority is violated (at institutions $h_{C}$), and $\nu_{*}(d_{Z})\succ_{d_{Z}}\beta(d_{Z})$. Additionally, note that $\nu_{*}$ Pareto-dominates $\beta$, i.e., no applicant $d$ exists with $\beta(d)\succ_{d}\nu_{*}(d)$ (regardless of whether the priority of $d$ is violated). Thus, $\nu_{*}$ is a priority-correcting adjustment of $\beta$, and $\beta\notin\mathsf{PN}$.
Next, consider the matchings $\alpha\beta$ and $\beta\gamma$:
$\alpha\beta:$ $h_{A}$ $h_{B}$ $h_{C}$ $h_{D}$ $h_{E}$ $h_{F}$ $h_{Z}$ $d_{B}$ $d_{A}$ $d_{D}$ $d_{C}$ $d_{E}$ $d_{F}$ $d_{Z}$     $\beta\gamma:$ $h_{A}$ $h_{B}$ $h_{C}$ $h_{D}$ $h_{E}$ $h_{F}$ $h_{Z}$ $d_{A}$ $d_{B}$ $d_{D}$ $d_{C}$ $d_{F}$ $d_{E}$ $d_{Z}$
We will show that both $\alpha\beta$ and $\beta\gamma$ are in $\mathsf{PN}$, i.e., we will show that these matchings have no priority-correcting adjustments. To begin to see this, observe directly that $d_{Z}$ is the only applicant whose priority is violated in these matchings, so any priority-correcting adjustment $\nu$ must improve the match of $d_{Z}$. Informally, we will show that no priority-correcting adjustment can exist because the only “reasonable” candidate will be $\nu_{*}$ as defined above. However, $\nu_{*}$ makes $d_{B}$ worse off than in $\alpha\beta$, and $d_{B}$’s priority is violated at $h_{A}$ in $\nu_{*}$, so $\nu_{*}$ is not a priority-correcting adjustment of $\alpha\beta$. Similarly, $\nu_{*}$ makes $d_{F}$ worse off than in $\beta\gamma$, and $d_{F}$’s priority is violated at $h_{E}$ in $\nu_{*}$, so $\nu_{*}$ is not a priority-correcting adjustment of $\beta\gamma$. Below, we show that indeed $\alpha\beta$ and $\beta\gamma$ have no priority-correcting adjustments.
We now formally show that $\alpha\beta\in\mathsf{PN}$. Suppose for contradiction that there exists a priority-correcting adjustment $\nu$ of $\alpha\beta$. Since the only applicant in $\alpha\beta$ whose priority is violated in $d_{Z}$, we know that $\nu$ must improve the match of $d_{Z}$, so $\nu(d_{Z})=h_{C}$ or $\nu(d_{Z})=h_{E}$. We consider these two cases. In the first case, if $\nu(d_{Z})=h_{C}$, then $\nu(d_{D})$ must be $h_{D}$ (since $d_{D}$ must receive a worse match in $\nu$ than in $\alpha\beta$, where $\alpha\beta(d_{D})=h_{C}$, and $d_{D}$ has highest-priority at $h_{D}$, which is $d_{D}$’s top choice after $h_{C}$). But, then $\nu(d_{C})$ must be $h_{C}$ (since similarly $d_{C}$ must receive a worse match in $\nu$ than in $\alpha\beta$, and $d_{C}$ has top priority at $h_{C}$), however, this contradicts the assumption that $d_{Z}$ is matched to $h_{C}$ in matching $\nu$. In the second case, suppose $\nu(d_{Z})=h_{E}$. Since $d_{E}$ has higher priority than $d_{Z}$ at $h_{E}$, we know $d_{E}$ must be matched in $\nu$ to an institution she prefers to $h_{E}$; so we either have $\nu(d_{E})=h_{A}$ or $\nu(d_{E})=h_{F}$. If $\nu(d_{E})=h_{A}$, then $d_{B}$ must receive a worse match in $\nu$ than in $\alpha\beta$; since $d_{B}$ has higher priority at $h_{A}$ than $d_{E}$, this would mean $\nu$ cannot be a priority-correcting adjustment, so we cannot have $\nu(h_{E})=h_{A}$. If $\nu(d_{E})=h_{F}$, then $\nu(d_{F})$ must be different from $\alpha\beta(d_{F})=h_{F}$; since $d_{F}$ has higher priority at $h_{E}$ than $d_{Z}$ and $d_{F}$’s preference list is $h_{E}\succ h_{F}$, this makes it impossible that $\nu(d_{Z})=h_{E}$, and we cannot have $\nu(d_{E})=h_{F}$. Thus, this shows that there exists no priority-correcting adjustment $\nu$ of $\alpha\beta$, so $\alpha\beta\in\mathsf{PN}$.
We now formally show that $\beta\gamma\in\mathsf{PN}$. Again, assume for contradiction there is a priority-correcting adjustment $\nu$ of $\beta\gamma$. Since $d_{Z}$ is the only applicant whose priority is violated in $\beta\gamma$, we must have $\nu(d_{Z})=h_{C}$ or $\nu(d_{Z})=h_{E}$. By precisely the same logic as in the previous paragraph, we cannot have $\nu(d_{Z})=h_{C}$ (since this case only considered the matches of $d_{C}$ and $d_{D}$, which are the same in $\beta\gamma$ as in $\alpha\beta$). Thus, consider $\nu$ such that $\nu(d_{Z})=h_{E}$. However, since $d_{F}$’s favorite institution was $h_{E}=\beta\gamma(d_{F})$, and $d_{F}$ has higher priority than $d_{Z}$ at $h_{E}$, we immediately get that $\nu$ is not a priority-correcting adjustment of $\beta\gamma$. Thus, this shows that there exists no priority-correcting adjustment $\nu$ of $\beta\gamma$, so $\beta\gamma\in\mathsf{PN}$.
Thus, the priority-neutral lattice consists of exactly those matchings in Figure 1, completing the proof. ∎
We can now use this example to give the main result of this paper. This resolves an open question of [Ren22]: the greatest lower bound of two matchings in $\mathsf{PN}$ need not be their coordinatewise minimum. It also shows that, crucially, the $\mathsf{PN}$ lattice is not distributive, and hence cannot be represented by a partial order on a set of rotations, as for the stable lattice.
Theorem 3.2 ($\mathsf{PN}$ need not be distributive).
$\mathsf{PN}$ may not be a distributive lattice. Moreover, the greatest lower bound of two matchings in $\mathsf{PN}$ may not be their coordinatewise minimum.
Proof.
We prove that both results of the theorem hold true in the lattice of priority-neutral matchings in Figure 1. To prove the second sentence, observe that
$\alpha\beta\wedge_{\mathsf{PN}}\beta\gamma=\mu_{0}\neq\beta=\alpha\beta\wedge_{\mathsf{L}}\beta\gamma,$
where we recall that $\wedge_{\mathsf{PN}}$ is the greatest lower bound of two matchings within $\mathsf{PN}$, and $\wedge_{\mathsf{L}}$ is their coordinatewise minimum. To prove the first sentence, we build on the above fact (used in equality (*)) and observe that
$\alpha\beta\wedge_{\mathsf{PN}}\big{(}\alpha\vee\beta\gamma\big{)}=\alpha\beta\wedge_{\mathsf{PN}}\alpha\beta\gamma=\alpha\beta\neq\alpha=\alpha\vee\mu_{0}\overset{(*)}{=}\big{(}\alpha\beta\wedge_{\mathsf{PN}}\alpha\big{)}\vee\big{(}\alpha\beta\wedge_{\mathsf{PN}}\beta\gamma\big{)}.$
Thus, the $\mathsf{PN}$ lattice is not distributive, which finishes the proof. ∎
4 Not All Lattices Arise From Priority-Neutral Matchings
In this section, we show that not all lattices arise as the set of priority-neutral matchings for some priorities and preferences. To prove this result, we give a novel structural property of the $\mathsf{PN}$ lattice, how it relates to the legal lattice $\mathsf{L}$ of [EM20], and about certain special types of chains of matchings, i.e. sequences $\mu_{1}\leq\mu_{2}\leq\ldots\leq\mu_{k}$. Then, we use these properties to rule out a specific lattice; as we discuss below, this lattice is in fact a classic example known as $M_{3}$ in the order theory literature.
4.1 Structural properties
4.1.1 Properties of $\mathsf{PN}$
To begin, we recall one of the integral tools [Ren22] uses to study $\mathsf{PN}$ matchings. This is the notion of a (Reny-)feasible chain, which is defined as follows:
Definition 4.1 ([Ren22]).
A chain $\mu_{1}<\mu_{2}<\ldots<\mu_{N}$ is Reny-feasible if $\mu_{1}$ is stable, $\mu_{N}$ is Pareto-efficient, and for each $i$, $\mu_{i}$ does not violate the priority of any applicant who is Pareto-improvable in $\mu_{i-1}$.
While [Ren22] reasons about Reny-feasible sequences in quite some detail, he does not prove the following fundamental fact about Reny-feasible sequences: every matching in a Reny-feasible sequence is always priority-neutral. To prove this fact, we extend the techniques from [Ren22, Lemmas 3–6]. First, we recall another simple lemma from [Ren22]:
Lemma 4.2 ([EM20], Lemmas 4–6; [Ren22], Lemma 1).
Let $\mu$ and $\nu$ be such that for all applicants $d$: $d$’s priority at $\mu(d)$ is not violated in $\nu$, and $d$’s priority at $\nu(d)$ is not violated in $\mu$. Then, $\mu\vee\nu$ and $\mu\wedge\nu$ are matchings.
We now prove our core result regarding $\mathsf{PN}$.
Proposition 4.3.
Let $\mu_{1}<\mu_{2}<\ldots<\mu_{N}$ be any Reny-feasible sequence. Then for each $n\in\{1,\ldots,N\}$, the matching $\mu_{n}$ is priority-neutral.
Proof.
Call $\xi$ an improver-only violator of $\mu$ if, for all $d\in\operatorname*{vio}(\xi)$, we have $\xi(d)\succeq_{d}\mu(d)$. Note that a priority-correcting adjustment of $\mu$ is simply an improver-only violator, which also improves the match of some applicant whose priority is violated by $\mu$. We proceed by considering any $n$ and any improver-only violator $\xi$ of $\mu_{n}$, and showing that $\xi$ cannot improve the match of any applicant whose priority is violated by $\mu_{n}$.
Fix some $n\in\{1,\ldots,N\}$, and suppose that $\xi$ is an improver-only violator of $\mu_{n}$. Using induction on $n^{\prime}$, we will prove the following claim: for all $n^{\prime}\in\{1,\ldots,n\}$, we have that for all applicants $d$,111111 [Ren22, Lemmas 4 and 5] perform related inductive arguments, but get weaker conclusions that are only sufficient to conclude that the final matching $\mu_{N}$ in the sequence is priority-neutral.
$d\in\operatorname*{vio}(\mu_{n^{\prime}})\qquad\implies\qquad\mu_{n^{\prime}}(d)\succeq_{d}\xi(d).$
The base case of $n^{\prime}=1$ follows from the fact that $\mu_{1}$ is stable by the definition of a Reny-feasible sequence, so $\operatorname*{vio}(\mu_{1})=\emptyset$.
Now assume by induction that the claim holds for some $n^{\prime}$. We will prove the claim for $n^{\prime}+1$.
To begin, we will prove using 4.2 that $\mu_{n^{\prime}}\vee\xi$ is a matching. To see this, observe that by induction, we know that for all $d\in\operatorname*{vio}(\mu_{n^{\prime}})$, we have $\mu_{n^{\prime}}(d)\succeq_{d}\xi(d)$. Thus, $\mu_{n^{\prime}}$ does not violate any $d$’s priority at $\xi(d)$. On the other hand, by the definition of an improver-only violator of $\mu_{n}$, we know that for all $d\in\operatorname*{vio}(\xi)$, we have $\xi(d)\succeq_{d}\mu_{n}(d)\succeq_{d}\mu_{n^{\prime}}(d)$. Thus, $\xi$ does not violate any $d$’s priority at $\mu_{n^{\prime}}(d)$. Thus, 4.2 implies that $\mu_{n^{\prime}}\vee\xi$ is a matching.
Now, consider any applicant $d\in\operatorname*{vio}(\mu_{n^{\prime}+1})$. By the definition of a Reny-feasible sequence, $d$ is not Pareto-improvable in $\mu_{n}^{\prime}$. Since $\mu_{n^{\prime}+1}\geq\mu_{n^{\prime}}$, also by the definition of a Reny-feasible sequence, we thus have $\mu_{n^{\prime}+1}(d)=\mu_{n^{\prime}}(d)$. Thus, since $\mu_{n^{\prime}}\vee\xi\geq\mu_{n^{\prime}}$, we must have $[\mu_{n^{\prime}}\vee\xi](d)=\mu_{n^{\prime}}(d)$, and therefore $\mu_{n^{\prime}+1}(d)=\mu_{n^{\prime}}(d)\succeq_{d}\xi(d)$. This finishes the proof of the inductive step, and proves the claim for all $n^{\prime}\leq n$.
This result directly implies that $\mu_{n}$ must be priority-neutral. To see this, we apply the claim with $n^{\prime}=n$. In detail, consider any $\xi$ which is a improver-only violator of $\mu_{n}$. By the claim, we know that for all $d\in\operatorname*{vio}(\mu_{n})$, we have $\mu_{n}(d)\succeq_{d}\xi(d)$. Thus, $\xi$ is not a priority-correcting adjustment of $\mu_{n}$. Thus, in fact $\mu_{n}$ has no priority-correcting adjustments, and in fact $\mu_{n}$ must be priority-neutral. ∎
4.1.2 Properties of $\mathsf{L}$
Now, we turn to the legal lattice $\mathsf{L}$. To begin, recall that [EM20] call a set of matchings $\mathsf{L}$ legal if $\mathsf{L}=\{\mu\ \allowbreak|\ \allowbreak\text{$\forall\nu\in\mathsf{L}:$ \allowbreak no applicant $d$'s priority is violated by $\mu$ at institution $\nu(d)$}\}$. [EM20] prove that there exists a unique legal set $\mathsf{L}$, and that legal matchings also form a lattice which contains all stable matchings, and whose applicant-optimal element equals $\mathsf{EADAM}$. Moreover, [Ren22] shows that $\mathsf{PN}\subseteq\mathsf{L}$.
Unlike priority-neutral lattices, however, legal lattices always have the same structural properties as stable lattices. [FZ22] formalizes this by showing that every legal lattice can be represented as the stable lattice on some altered set of preferences and priorities; we make heavy use of this idea, and other techniques from [FZ22], below. In summary, the main known results are the following:
Theorem 4.4 ([EM20, Ren22, FZ22]).
We have $\mathsf{PN}\subseteq\mathsf{L}$, and $\mathsf{L}$ forms a distributive lattice which can be represented via a partial order on a set of rotations over some other set of preferences and priorities.
Now, we introduce another type of chain in order to additionally reason about the legal lattice while bringing in the techniques of [TY14]. This chain modifies the Kesten-Tang-Yu chain (as defined in 2.3) by, informally speaking, progressing the chain with “as small of differences as possible.” To formalize this, we recall that in any lattice, an element $x$ covers $y$ if $x\geq y$, and for any $x\geq z\geq y$, we have either $x=z$ or $y=z$. Covering elements in stable lattice correspond to two matchings differing by rotations in the rotation poset. In other words, $\mu_{2}$ covers $\mu_{1}$ in a stable lattice if and only if there is a rotation $\alpha$ exposed in $\mu_{1}$ such that $\alpha\mu_{1}=\mu_{2}$ (by the notation of Section 2.1). See [GI89] (where covering relationships are discussed under the term “minimal differences”) or [CT19] for details.
Definition 4.5.
We say that a chain of matchings is a gradual Kesten-Tang-Yu chain if it can be constructed as follows: Set $P_{0}=P$ to be the full preference (and priority) profile, and let $\mu_{0}=\mathsf{IPDA}(P_{0})$. Now, for $i\geq 1$, as long as $\mu_{i-1}\neq\mathsf{APDA}(P_{i-1})$, set $P_{i}=P_{i-1}$, and set $\mu_{i}$ equal to any matching which covers $\mu_{i-1}$ in the stable matching lattice under preferences $P_{i-1}$. If $\mu_{i-1}=\mathsf{APDA}(P_{i-1})$, yet $\mu_{i-1}$ is not Pareto-optimal, then modify $P_{i}$ as in 2.3 (i.e., let $P_{i}$ denote the following operation on $P_{i-1}$: for each $d$ matched to an under-demanded institution in $\mu_{i-1}$, remove all institutions in their list strictly before $\mu_{i-1}(d)$). One can show $\mu_{i-1}$ is stable under preferences $P_{i}$ as well; we set $\mu_{i}$ to any matching stable under $P_{i}$ which covers $\mu_{i-1}$ in the stable lattice under $P_{i}$. The chain terminates with $\mu_{N}$ once $\mu_{N}$ is Pareto-optimal.
By [Ren22, Lemma 2], or by the techniques of [TY14], we know that if $\mu_{i}$ in the gradual Kesten-Tang-Yu chain is not Pareto optimal, then some applicants must be matched to under-demanded institutions.121212 In more detail, [Ren22, Lemma 2] shows that if an applicant $d$ is Pareto-improvable at $\mu_{i}=\mathsf{APDA}(P_{i})$, then $d$ cannot be matched to an under-demanded in $\mathsf{APDA}(P_{i})$. Thus, whenever $\mu_{i}=\mathsf{APDA}(P_{i})$ but $\mu_{i}$ is not Pareto-optimal, we know that $P_{i+1}\neq P_{i}$. Conversely, if the matching is Pareto-optimal, then it must equal $\mathsf{APDA}$ under the current preferences (and under the original preferences). Hence there always exists at least one gradual Kesten-Tang-Yu sequence.
Additionally, any gradual Kesten-Tang-Yu sequence is also Reny-feasible. To see why, observe that, by construction, a gradual Kesten-Tang-Yu sequence will only violated the priority of applicants matched to under-demanded institutions. Moreover, the first matching $\mu_{0}$ of a gradual Kesten-Tang-Yu chain equals $\mathsf{IPDA}$, and the last matching $\mu_{N}$ is Pareto-efficient, by definition. Hence, by 4.3, all elements of any gradual Kesten-Tang-Yu sequence are priority-neutral, and by Theorem 4.4, also legal.
We now use results from [FZ22] and [TY14] to establish that any gradual Kesten-Tang-Yu sequence consists of covering relationships in the legal lattice; i.e., any gradual Kesten-Tang-Yu sequence is a chain in the legal lattice which starts with $\mathsf{IPDA}$, and is maximal among such chains. This constitutes our core result regarding $\mathsf{L}$.
Proposition 4.6.
Let $\mu_{1}<\mu_{2}<\ldots<\mu_{N}$ be any gradual Kesten-Tang-Yu sequence. Then, for each $n=1,\ldots,N-1$, we have that $\mu_{n}$ is covered by $\mu_{n+1}$ in $\mathsf{L}$.
Proof.
Fix some profile $P$ of preferences (and priorities). To begin, we recall some concepts from [FZ22]. If applicant $d$ is never matched to institution $h$ in any matching in $\mathsf{L}$, then [FZ22] call the pair $(d,h)$ an illegal pair. Let $\widetilde{P}$ denotes the altered preferences where, for all $d$, we modify $d$’s preference list by removing all $h$ such that $(d,h)$ is an illegal pair. [FZ22, Lemma 3.5] shows that the legal lattice on $P$ equals the stable lattice on $\widetilde{P}$ (and moreover that illegal pairs can always be removed from applicants preference lists without changing the set of legal matchings). Our proof proceeds by showing that $\widetilde{P}$ can be constructed from the execution of a gradual Kesten-Tang-Yu algorithm as in 4.5 (along with additional steps after the sequence) in a way such that the successive elements of the gradual Kesten-Tang-Yu sequence correspond covering relationships in the stable lattice under $\widetilde{P}$.
Now, consider $\mu_{0}=\mathsf{IPDA}(P)$. We know from the main result of [TY14] (or [Ren22, Lemma 2]) that in 4.5, every $d$ matched to an under-demanded institution must be matched as in $\mathsf{EADAM}$. Thus, for each $h^{\prime}$ above such a $d$’s current match in $\mu_{i}$ is such that $(d,h^{\prime})$ is an illegal pair. Thus, for each step of 4.5, for each $d$ matched to an under-demanded institution, all of the institutions preferred to $d$’s current match $h$ can be removed from $d$’s list without changing the set of legal matchings. Let $P_{\mathsf{top}}$ denote the result of applying this operation successively for all steps of 4.5. Note that there are still additional illegal edges that remain on applicants’ lists in $P_{\mathsf{top}}$; since we know all matchings $\mu_{i}$ are legal (by 4.3 and the fact that $\mathsf{PN}\subseteq\mathsf{L}$), these remaining illegal pairs are some set of pairs $(d,h)$ such that $\mu_{i}(d)\prec_{d}h\prec_{d}\mu_{i+1}(d)$ for some $i$ (or such that $h\prec_{d}\mu_{1}=\mathsf{IPDA}$). However, we can see from the definition of rotations under some set of preferences (see e.g. [GI89]) that deleting any such $h$ from $d$’s list will never change whether a given rotation $\alpha$ is an actual rotation under the preferences; in the language of [GI89], a rotation which was exposed in a matching $\mu$ before any further removal of such illegal pairs must remail exposed in $\mu$ after we remove these further illegal pairs, directly by the definition of a rotation exposed in $\mu$. (Note, however, that this operation may change the predecessor relationships between the rotations, to see this, consider e.g. and type II predecessor relationships as defined in [GI89]—this predecessor relationship will be removed in the legal rotation poset.) Thus, when all illegal edges have been removed, and thus by [FZ22, Lemma 3.6] the legal set equals the stable set, we will still have that all rotations along the gradual Kesten-Tang-Yu sequence also constitute rotations in the legal lattice. Since one matching covers another in a stable matching lattice if and only if they differ by rotations, the result follows. ∎
4.1.3 Structural result
We now prove our structural result for $\mathsf{PN}$. To state this structural result, we again make heavy use of the notion of the rotation poset [GI89]. First, we introduce some notation.
Let $\mathsf{K}$ denote the sub-lattice of the legal lattice $\mathsf{L}$ which consists of all matchings that dominate $\mathsf{IPDA}$, i.e., $\mathsf{K}=\{\mu\in\mathsf{L}\ \mid\ \mu\geq\mathsf{IPDA}\}$.131313 Lattice $\mathsf{K}$ is more convenient for our purposes than $\mathsf{L}$, because in addition to maintaining the same nice structure as $\mathsf{L}$, lattice $\mathsf{K}$ has the same highest and lowest elements as $\mathsf{PN}$. This means that the lowest element of $\mathsf{PN}$, namely $\mathsf{IPDA}$ (Theorem 2.5), corresponds to the empty set of rotations in the representation of $\mathsf{K}$ in Theorem 4.7. (In contrast, the legal lattice $\mathsf{L}$ always contains the institution-side analogue of $\mathsf{EADAM}$—i.e., the matching resulting from running $\mathsf{EADAM}$ while interchanging the roles of applicants and institutions—because the definition of $\mathsf{L}$ is symmetric in the roles of applicants and institutions. Hence $\mathsf{L}$ can have a bottommost element strictly below $\mathsf{IPDA}$.) Note that $\mathsf{PN}\subseteq\mathsf{K}$ by Theorem 2.5 and Theorem 4.4. From the fact that $\mathsf{L}$ is a distributive lattice, we immediately see that $\mathsf{K}$ is a distributive lattice; hence it can be represented by a partial order on some set. We will moreover show as a simple corollary of [FZ22] a that $\mathsf{K}$ can be represented by a partial order on a set of rotations that one applies to some base matching $\mu_{0}$, exactly as for the set of stable matchings and as for $\mathsf{L}$. Our main structural result will combine 4.3 and 4.6 in order to additionally show that there is some maximal chain in $\mathsf{K}$ such that every element in this maximal chain is also in $\mathsf{PN}$.
Theorem 4.7 (Structural Result for $\mathsf{PN}$).
Fix any profile of preferences and priorities. There exists a set of rotations $R$, and a partial order on those rotations, such that $\mathsf{K}\supseteq\mathsf{PN}$ is isomorphic to the downward-closed subsets of the partial order $R$.
Moreover, there exists an ordering $[\alpha_{1},\ldots,\alpha_{N}]$ of the elements of $R$ such that for each $k\in\{1,\ldots,N\}$, we have $\alpha_{k}\alpha_{k-1}\ldots\alpha_{1}\mathsf{IPDA}\in\mathsf{PN}$.
Proof.
[FZ22] construct such a partial order $R^{\prime}$ for the lattice $\mathsf{L}$. To convert this partial order $R$ to work for $\mathsf{K}$, simply let $S$ be the (downward-closed) subset $R^{\prime}$ corresponding to $\mathsf{IPDA}$, and remove all rotations in $S$ to produce $R$. Then, it is straightforward to show that $\mathsf{K}$ is in a bijection with the downward closed subsets of $R$, with the isomorphism given (as always, see e.g. [GI89]) by mapping a downward-closed set $S=\{\alpha_{1},\ldots,\alpha_{k}\}$ to $\alpha_{k}\alpha_{k-1}\ldots\alpha_{1}\mathsf{IPDA}$ (possibly reordering the $\alpha_{i}$ such that $\alpha_{x}$ is never below $\alpha_{y}$ in the partial order for $x\geq y$). This proves the first paragraph of the theorem.
Now, consider any gradual Kesten-Tang-Yu sequence $\mathsf{IPDA}=\mu_{1}<\mu_{2}<\ldots<\mu_{N}=\mathsf{EADAM}$. Since each $\mu_{i}\in\mathsf{K}$, they correspond to a set $S_{i}\subseteq R$; also note that $S_{1}=\emptyset$. Moreover, by 4.6, each $\mu_{i}$ covers $\mu_{i-1}$ in $\mathsf{L}$, and hence in $\mathsf{K}$. Thus, for each $i=1,\ldots,N$, we have $S_{i}=S_{i-1}\cup\{\alpha_{i}\}$ for some $\alpha_{i}\in R$ (since covering relationships in lattices arising from rotation posets always correspond to adding a single rotation). Additionally, by 4.3, each $\mu_{i}\in\mathsf{PN}$. So, with respect to the ordering $[\alpha_{1},\alpha_{2},\ldots,\alpha_{N}]$ constructed as above, we know that for each $k\in\{1,\ldots,N\}$, we have $\alpha_{k}\alpha_{k-1}\ldots\alpha_{1}\mathsf{IPDA}\in\mathsf{PN}$. This proves the second paragraph of the theorem. ∎
4.2 $M_{3}$ is not a priority-neutral lattice
Figure 2 displays a lattice, typically denoted by $M_{3}$, which is a famous simple example in the lattice theory literature [Bir40, Grä12]. We use the facts proven in Section 4.1 to prove that $M_{3}$ never arises as the lattice of priority-neutral matchings.
(a) The lattice $M_{3}$.
(b) Key steps of the proof.
Figure 2: $M_{3}$ lattice and key steps of the proof of Theorem 4.8.
Theorem 4.8 ($M_{3}$ is not a $\mathsf{PN}$ lattice).
There is no profile of priorities and preferences such that $\mathsf{PN}$ is isomorphic to the lattice $M_{3}$ shown in Figure 2.
Proof.
Suppose for contradiction there is some set of preferences by which $\mathsf{PN}$ is isomorphic to $M_{3}$.
Using Theorem 4.7, consider the partial order $R$ on rotations which represents $\mathsf{K}$ (as defined above), and let $[\alpha_{1},\alpha_{2},\ldots,\alpha_{N}]$ be as in Theorem 4.7. Observe that this means there is some element $\mu_{1}\in\mathsf{PN}$ which corresponds to $\{\alpha_{1}\}\subseteq R$.
Now, since we have assumed $\mathsf{PN}$ is isomorphic to $M_{3}$, there are two other matchings, call them $\mu_{s}\neq\mu_{t}$, which cover $\mathsf{IPDA}$ in $\mathsf{PN}$. Suppose these matchings correspond to sets of rotations $S\subseteq R$ and $T\subseteq R$. Now, observe that since $\mathsf{PN}$ is closed under coordinatewise maxima (Theorem 2.5), we have that the join of two elements of $\mathsf{PN}$ corresponds to the union of the sets of rotations corresponding to the two elements. Thus, $\mu_{1}\vee_{\mathsf{PN}}\mu_{s}$ corresponds to $\{\alpha_{1}\}\cup S$, and $\mu_{1}\vee_{\mathsf{PN}}\mu_{t}$ corresponds to $\{\alpha_{1}\}\cup T$.
However, by the definition of $M_{3}$, we know that $\mu_{1}\vee_{\mathsf{PN}}\mu_{s}=\mu_{1}\vee_{\mathsf{PN}}\mu_{t}$. Thus, we must have $\{\alpha\}\cup S=\{\alpha\}\cup T$. (For an illustration of this argument, see Figure 2.) This implies that $S=T$141414 This step of the proof would fail if we were not able to establish that at least one matching in $\mathsf{PN}$ differed from $\mathsf{IPDA}$ by only one rotation. Indeed, if we replaced $\{\alpha_{1}\}$ with a set of just two rotations $\{\alpha_{1},\beta\}$, then we would not be able to conclude from $\{\alpha_{1},\beta\}\cup S=\{\alpha_{1},\beta\}\cup T$ that $S=T$. This shows how we use our structural result Theorem 4.7. But this contradicts the assumption that $S$ and $T$ correspond to distinct elements of the $\mathsf{PN}$ lattice. This concludes the proof. ∎
In closing, we make one final remark. The order theory literature calls a lattice modular if $a\leq b$ implies that for all $x$, we have $a\vee(x\wedge b)=(a\vee x)\wedge b$. One can show that every distributive lattice is modular, and indeed modular lattices form a canonical well-studied class of lattices capturing many structures [Bir40, Grä12]. The lattice $M_{3}$ in Figure 2 is a classic example of a modular but not distributive lattice. Thus, Theorem 4.8 shows that not all modular lattices arise as priority-neutral lattices. On the other hand, the $\mathsf{PN}$ lattice in Figure 1 is not modular, since $\alpha\vee(\beta\gamma\wedge_{\mathsf{PN}}\alpha\beta)=\alpha\neq\alpha\beta=(\alpha\vee\beta\gamma)\wedge_{\mathsf{PN}}\alpha\beta$. Thus, not all priority-neutral lattices are modular. Thus, the class of modular lattices—a canonical generalization of distributive lattices—is unrelated (in the set-inclusion sense) to the class of $\mathsf{PN}$ lattices. Future work may investigate additional order-theoretic properties and hope to characterize the class of priority-neutral lattices, or may find further complex and intractable aspects to priority-neutral lattices.